Your company is buying a new piece of equipment to reduce emissions in your factory. It will cost​ $2,000 to run next year​ (Year 1), and that operating cost will increase by​ 4% every year until Year​ 20, at which point you will discard it​ (after paying the operating cost​ - there is a cash flow in Year​ 20). You will also have to pay to maintain the​ equipment, which you estimate will cost​ $400 next year​ (Year 1) and increase by​ 5% per year until Year 20. Assume an interest rate of​ 9%.

a. What is the equivalent present value in Year 0 of the total costs of the​ equipment?

b. What is the equivalent future value in Year 20 of the total costs of the​ equipment?   
Answer #1

Answer a. Let's calculate present value (PV) of operating cost and maintainance cost. PV = C / (1+r)^n
Where, c is cashflow in a year, r is rate of interest and n is time period.
Note that operating expense in first year is $2000 and increases by 4% every year. Maintainance cost is $400 and increases by 5% per year. Interest rate is 9%. Following table shows operating and maintainance expenses for entire time period in column 1 and column 2 respectively. column 3 and column 4 shows the present value of operating and maintainance expenses respectively. Sum of column 3 and column 4 will give equivallent present value of total cost of equipment in year 0.

Year Operating cost (OC) Maintainance cost (MC) PV of OC PV of MC
1 2000 400 1834.86 366.97
2 2080 420 1750.69 353.51
3 2163.2 441 1670.39 340.53
4 2249.73 463.05 1593.76 328.04
5 2339.72 486.20 1520.66 316.00
6 2433.31 510.51 1450.90 304.40
7 2530.64 536.04 1384.35 293.23
8 2631.86 562.84 1320.84 282.47
9 2737.14 590.98 1260.25 272.10
10 2846.62 620.53 1202.44 262.12
11 2960.49 651.56 1147.29 252.50
12 3078.91 684.14 1094.66 243.23
13 3202.06 718.34 1044.45 234.31
14 3330.15 754.26 996.53 225.71
15 3463.35 791.97 950.82 217.43
16 3601.89 831.57 907.21 209.45
17 3745.96 873.15 865.59 201.76
18 3895.80 916.81 825.89 194.36
19 4051.63 962.65 788.00 187.23
20 4213.70 1010.78 751.85 180.35

Total equivallent present cost = 29627.14 (sum of all present values calculated in column 3 and column 4).

Answer B. Let's now calculate future value of total cost of equipment. Future value = amount * (1+r)^n

Where, r is rate of interest and n is time period.

Here, amount is $29627.14 (from answer A).

Future value = 29627.14 * 1.09^20 = 166042.6.